OpenStudy (anonymous):

Solve for x: 3x − 5 = 2x + 6

sam (.sam.):

\[\begin{array}{l} 3 x-5=2 x+6 \\ \text{Add }(5-2 x)\text{\to both sides} \\ x=11 \\\end{array}\]

OpenStudy (anonymous):

3x − 5 = 2x + 6 //add 5 3x - 5 + 5 = 2x + 6 + 5 3x = 2x +11 //subtract 2x 3x-2x = 2x-2x +11 x = 11